OpenStudy (anonymous):

i invest $1000 in an account earning 6% compounded annually. how long will it take for the value to reach $1500? Im close but not sure..

OpenStudy (anonymous):

\[1000\times1.06^t=1500=>1.06^t=3/2=>t=\log_{1.06}3/2=6.95\]So it takes seven years.